# D-inverse constellations

### v1

@@ 6. Inverse constellations via pre-constellations
The definition of inverse semigroups does not involve domain or range operations. We conclude this article by giving a formulation of the inverse constellation concept that does not presuppose the existence of a domain operation.
@@
[…]
As usual, we may sometimes denote a pre-constellation $(P,\cdot)$ simply by “$P$”. Every semigroup is nothing but a pre-constellation in which all products are defined.
In a constellation, $RI(\mathcal{P})=D(\mathcal{P})\subseteq E(\mathcal{P})$ as seen in Section 2. It follows that $D(\mathcal{P})$ is determined by the structure of $\mathcal{P}$ as a constellation. Hence, a pre-constellation is a constellation in at most one way. In general, $D(\mathcal{P})$ and $E(\mathcal{P})$ can be different, as may easily be seen by considering $\mathcal{C}_{X}$. However, by Proposition 3.2, if $\mathcal{P}$ is inverse then $RI(\mathcal{P})=D(\mathcal{P})=E(\mathcal{P})$.
We wish to define a notion of regularity in a pre-constellation. As at the beginning of Subsection 3.1, a semigroup is regular if for all $a\in S$ there is a $b\in S$ such that $a=aba$. By setting $c=bab$ it then follows that for all $a$ there is $c$ such that $a=aca$ and $c=cac$.
In the case of (pre-)constellations, there is ambiguity in the definition of $x\cdot y\cdot x$, so we must take a little care.
@@
and, continuing to make use of (Const1) and (Const2)
@@
If $\mathcal{P}$ is a pre-constellation, we say that $t\in P$ is an inverse of $s\in P$ if $s\cdot(t\cdot s)=s$ and $t=t\cdot(s\cdot t)$, and $P$ is an inverse pre-constellation if every element of $P$ has a unique inverse.
@@
Suppose $\mathcal{P}$ is a regular pre-constellation. Then it is inverse if and only if for all $e,f\in E(\mathcal{P})$, if $e=e\cdot(f\cdot e)$ and $f=f\cdot(e\cdot f)$, then $e=f$.
@@
Conversely, suppose $P$ is regular and the stated condition on idempotents holds. Suppose $s\in P$ has both $t,u\in P$ as inverses. So $s=s\cdot(t\cdot s)$, $t=t\cdot(s\cdot t)$, $s=s\cdot(u\cdot s)$ and $u=u\cdot(s\cdot u)$. It follows that $s\cdot t,s\cdot u,t\cdot s$ and $u\cdot s$ are all idempotent. One can then calculate in a standard way that
@@
If $P$ is a semigroup, the last result asserts that if $\mathcal{P}$ is regular then it is inverse if and only if whenever $e=efe$ and $f=fef$ for two idempotents $e,f$, it must be that $e=f$, a condition which is therefore equivalent to the commuting idempotents condition most frequently used.
@@
Every D-inverse inverse constellation $\mathcal{P}=(P,\cdot\,,D)$ has an inverse pre-constellation reduct $(P,\cdot)$ in which the inverses of elements are their D-inverses, satisfying the following:
@@
Suppose $\mathcal{P}$ is a D-inverse constellation. For all $s\in P$ we have $s\cdot(s^{\prime}\cdot s)=s\cdot R(s)=s$, and similarly s($s^{\prime}\cdot(s\cdot s^{\prime})=s^{\prime}$, so that $\mathcal{P}$ is regular.
@@
The two conditions in Proposition 6.6 characterise (reducts of) D-inverse constellations amongst inverse pre-constellations.
@@
If $\mathcal{P}$ is a pre-constellation that is inverse and satisfies the two conditions in Proposition 6.6, then it is a D-inverse constellation in which the inverse $s^{\prime}$ of $s\in P$ is its D-inverse (and hence $D(s)=s\cdot s^{\prime}$ for all $s\in P$).

### v2

@@ 6. Inverse constellations via pre-constellations
The definition of inverse semigroups does not involve domain or range operations. We conclude by giving a formulation of the D-inverse constellation concept that does not presuppose the existence of a domain operation.
@@
We say that $\mathcal{P}=(P,\cdot)$ is a pre-constellation if $\cdot$ is a partial binary operation that satisfies ((Const1)) and ((Const2)) in the definition of a constellation.
Every semigroup is nothing but a pre-constellation in which all products are defined.
In a constellation, $RI(\mathcal{P})=D(\mathcal{P})\subseteq E(\mathcal{P})$ as seen in Section 2. It follows that $D(\mathcal{P})$ is determined by the structure of $\mathcal{P}$ as a constellation. Hence, a pre-constellation is a constellation in at most one way. In general, $D(\mathcal{P})$ and $E(\mathcal{P})$ can be different, as may be easily seen by considering $\mathcal{C}_{X}$. However, by Proposition 3.2, if $\mathcal{P}$ is inverse then $RI(\mathcal{P})=D(\mathcal{P})=E(\mathcal{P})$.
We wish to define a notion of regularity in a pre-constellation. Recall from Section 3 that a semigroup is regular if for all $a\in S$ there exists a $b\in S$ such that $a=aba$; in this case, by setting $c=bab$ (where $a=aba$), it follows that $a=aca$ and $c=cac$.
In the case of pre-constellations, there is ambiguity in the definition of $x\cdot y\cdot x$, so we must take a little care.
@@
and, continuing to make use of (Const1) and (Const2),
@@
If $\mathcal{P}$ is a pre-constellation, we say that $t\in P$ is an inverse of $s\in P$ if $s\cdot(t\cdot s)=s$ and $t=t\cdot(s\cdot t)$, and $\mathcal{P}$ is an inverse pre-constellation if every element of $\mathcal{P}$ has a unique inverse.
@@
Suppose that $\mathcal{P}$ is a regular pre-constellation. Then it is inverse if and only if for all $e,f\in E(\mathcal{P})$, if $e=e\cdot(f\cdot e)$ and $f=f\cdot(e\cdot f)$, then $e=f$.
@@
Conversely, suppose that $\mathcal{P}$ is regular and the stated condition on idempotents holds. Suppose that $s\in P$ has both $t,u\in P$ as inverses. So $s=s\cdot(t\cdot s)$, $t=t\cdot(s\cdot t)$, $s=s\cdot(u\cdot s)$ and $u=u\cdot(s\cdot u)$. It follows that $s\cdot t,s\cdot u,t\cdot s$ and $u\cdot s$ are all idempotent. One can then calculate in a standard way that
@@
If $\mathcal{P}$ is a semigroup, the last result asserts that if $\mathcal{P}$ is regular then it is inverse if and only if whenever $e=efe$ and $f=fef$ for two idempotents $e,f$, it must be that $e=f$, a condition which is therefore equivalent to the more familiar commuting idempotents characterisation.
@@
Every D-inverse constellation $\mathcal{P}=(P,\cdot\,,D)$ has an inverse pre-constellation reduct $(P,\cdot)$ in which the inverses of elements are their D-inverses, satisfying the following:
@@
Suppose that $\mathcal{P}$ is a D-inverse constellation. For all $s\in P$ we have $s\cdot(s^{\prime}\cdot s)=s\cdot R(s)=s$, and similarly $s^{\prime}\cdot(s\cdot s^{\prime})=s^{\prime}$, so that $\mathcal{P}$ is regular.
@@
Next we show that the two conditions in Proposition 6.6 in fact characterise (reducts of) D-inverse constellations amongst inverse pre-constellations.
@@
If $\mathcal{P}$ is a pre-constellation that is inverse and satisfies the two conditions in Proposition 6.6, then it is a D-inverse constellation in which the inverse $s^{\prime}$ of $s\in P$ is its D-inverse ((and hence $D(s)=s\cdot s^{\prime}$ for all $s\in P$).
@@
Acknowledgement
The authors are grateful to a careful referee, whose remarks have helped improve our paper.